OpenStudy (anonymous):

Find the volumes of the solids generated by revolving the regions bounded by the graphs of the equations about the given lines. y = sqrt(x) y = 0 x = 6 a) the x-axis b) the y-axis c) the line x=6 d) the line x=9

OpenStudy (anonymous):

I'll start you off:  a) Use the method of rings/washers, \[r = \sqrt{x}\] and integrate from x=0 to x=6 b) Use the method of cylindrical shells. \[h = \sqrt{x}\] and integrate from x=0 to x=6 c) Same as b except think about where you're rotating, how will your radius change? d) Once you've done c, you can do this For more information see: http://tutorial.math.lamar.edu/Classes/CalcI/VolumeWithCylinder.aspx

OpenStudy (anonymous):

Its x^(1/2) in the picture not x^2

OpenStudy (anonymous):

So for a) would it be \[\int\limits_{0}^{6}x ^{1/2} dx\]? If so, how would b) be any different? I still dont understand C and D